OpenStudy (anonymous):

Simplify the expression: Sin^2x-1/cos(-x)

OpenStudy (solomonzelman):

Have you seen the following identity? \(\large\color{black}{ \displaystyle \sin^2\theta+\cos^2\theta=1 }\)

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

Subtract \(\sin^2\theta\) from both sides and tell me what you get.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \sin^2\theta+\cos^2\theta =1 \ }\) \(\large\color{black}{ \displaystyle \sin^2\theta+\cos^2\theta \color{blue}{-\sin^2\theta }=1 \color{blue}{-\sin^2\theta } }\) \(\large\color{black}{ \displaystyle \cos^2\theta =1-\sin^2\theta }\)

OpenStudy (solomonzelman):

So, based on the very last statement, you can now simplify your expression: \(\large\color{black}{ \displaystyle \frac{1-\sin^2x}{\cos x} =\frac{?}{\cos x} }\)

OpenStudy (solomonzelman):

hello?

OpenStudy (anonymous):

Sinx?

OpenStudy (solomonzelman):

We have concluded that: \(1-\sin^2\theta =\cos^2\theta\) correct?

OpenStudy (anonymous):

Yes so the ? Represents cos^2 theta?

OpenStudy (solomonzelman):

Yes, therefore it comes out that: \(\large\color{black}{ \displaystyle \frac{1-\sin^2x}{\cos x} =\frac{\cos^2x}{\cos x} }\)

OpenStudy (solomonzelman):

and THAT \(\Uparrow\) you can probably tell me how to simplify:D

OpenStudy (anonymous):

So the answer is cos x? :)

OpenStudy (solomonzelman):

yes, just \(\cos x\).

OpenStudy (solomonzelman):

Any questions?

OpenStudy (anonymous):

Thank you for your help! i understand it now :D

OpenStudy (solomonzelman):

Yes, one more note: \(\color{red}{\cos(-\theta)=\cos(\theta)}\)

OpenStudy (solomonzelman):

So, lets quicly recap everything once more, okay?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{1-\sin^2(x)}{\cos (-x)} = }\) \(\large\color{black}{ \displaystyle \frac{1-\sin^2(x)}{\cos (x)} = }\) \(\large\color{black}{ \displaystyle \frac{\cos^2(x)}{\cos (x)} = \cos(x) }\)